A car's value depreciates at a rate of 20%

each year. If you paid $20,000 for this car 6

years ago, what is the current value of your

car?

y = a(1 - r)^t


1
Expert's answer
2021-03-25T06:44:44-0400

Given that

a=20,000

r=20%

t=6yrs

Using

"y = a(1-r)^t"

"y = 20000(1- \frac{20}{100})^6 = 20000(0.8)^6 = 2×10^4 × 8^6×10^{-6} = 5242.88"

The current value of the car is $5,242.88